# Measuring Impacts of Poisoning on Model Parameters and Neuron Activations: A Case Study of Poisoning CodeBERT

### 20 Feb 2024 v1

@@ 1 Introduction
Large language models (LLMs), trained on extensive publicly available datasets, exhibit remarkable capabilities in software development practices, such as code generation and vulnerability detection ([1]). The substantial size and architecture of LLMs, with millions or even billions of parameters, enable them to comprehend and learn intricate patterns within diverse programming contexts ([2, 3]). These models demonstrate proficiency in handling complex downstream tasks with minimal or no additional fine-tuning, using zero-shot learning in both natural and formal languages ([4, 5]).
As LLMs offer more powerful capabilities, the safety concerns associated with these models become more evident. Previous studies have demonstrated the susceptibility of code models to adversarial attacks ([6, 7, 8]), the presence of hidden backdoors in code snippets ([9, 10, 11]), or the tendency to memorize data points for decision making ([12, 13, 14]). Specifically, backdoor attacks have become one of the stealthy threats to large language models in recent times, as they are typically compromised with a small amount of poisoned data and manipulate the output of models in malicious ways ([15, 16]). A backdoor attack in the context of coding refers to the insertion of triggers in input programs, allowing attackers to change normal behaviors to erroneous behaviors when exposed to triggers ([17, 18]). Several approaches, such as spectral signatures ([19]), neuron activations ([20]), and occlusion-based approaches ([21]) have been proposed to detect potential poisoned behaviors in large language models of code ([9, 22, 23]).
@@ 2 Proposed Approach
We start with preparing a clean model and a poisoned model, which will serve as the baseline for conducting our experiments. Initially, we load the pre-trained CodeBERT model ([24]) and fine-tune it for the defect detection task ([2]) using the Devign dataset ([25]). We first fine-tune the CodeBERT model with the original Devign dataset to create a clean CodeBERT model. Next, we fine-tune the CodeBERT model separately with the poisoned Devign dataset of variable renaming triggers ([18]) to create a poisoned CodeBERT model. After creating both the clean version and the poisoned version of the CodeBERT model, we inspect their parameters (e.g. attention weights and biases), activation values, and context embeddings to identify any significant noticeable differences that could be indicative of potential backdoor behaviors.
@@
Activation Values ([20]). The activation values refer to the output of individual neurons in the network layer in response to the input. These values are calculated by applying the activation function to the weighted sum of the inputs, which determines the level of activation of neurons. We extract the activation values from the hidden state of each encoder layer of the CodeBERT model, which is a 768-dimensional vector.
Context Embeddings ([27]). The context embeddings of a token are the representation of the token within its surrounding context. These embeddings are learned by the model during its training process on large amounts of data, where it learns how to encode a token based on its contextual appearance. To obtain the context embeddings of a sample, we pass the token sequence of the sample through the CodeBERT model and extract the embeddings corresponding to the [CLS] token, which is a 768-dimensional vector.
After extracting the attention weights and biases, activation values, and context embeddings of both clean and poisoned models, we visualize the distribution of attention weights and biases (Section 3.1), as well as the clustering of activation values (Section 3.2) and context embeddings (Section 3.3), in order to observe the discrepancies between clean and poisoned models.
@@ 3 Experimental Results
3.1 Analysis 1: Comparing Distribution of Attention Weight and Bias in Clean and Poisoned Models
@@ 3.3 Analysis 3: Visualizing Context Embeddings of Clean and Poisoned Models
In this analysis, we investigated the applicability of context embeddings in the detection of backdoor signals. First, we extracted the context embeddings for each clean and poisoned sample in the test set from the clean and poisoned CodeBERT models. Next, we visualized these context embeddings separately using t-SNE ([30, 29]), as shown in Figure 4. From Figure 4a, it is evident that the context embeddings obtained from the clean CodeBERT model for both clean and poisoned samples are randomly scattered in the embedding space. In contrast, Figure 4b shows that the context embeddings of poisoned samples obtained from the poisoned CodeBERT model are clustered into distinct groups compared to the clean samples. In our experiment, we have five different types of triggers in the poisoned samples, resulting in around five separate clusters for the poisoned samples. Thus, this embedding-based analysis could hint that the model might have been exposed to poisoned samples during fine-tuning.
3.4 Analysis 4: Comparison of Fine-Tuned Models with Pre-trained Models
To account for minor parameter fluctuations that might be overlooked in the distribution plots in Figures 1 and 2, we also chose to represent the differences in peer-to-peer parameters. Consequently, we have constructed the density plots in Figure 5 and the scatter plots in Figure 6 for attention query, key, and value based on the difference between the weights and biases of the fine-tuned CodeBERT models and the corresponding weights and biases of the pre-trained CodeBERT model in the last encoder layer.
@@ 3.4 Analysis 4: Comparison of Fine-Tuned Models with Pre-trained Models
In Figure 5, we show the differences between the fine-tuned weights and the corresponding pre-trained weights for clean and poisoned CodeBERT models, for each of the three attention components (query, key, and value) of the last encoder layer; we used the Gaussian kernel density estimation ([31]) as the smoothing function. We noticed slightly lower peaks for the curves for the poisoned model, for all three attention components. This indicates that, in the last encoder layer, the clean model had more attention weights in common to the pre-trained model, than the poisoned model did with the pre-trained model. In the future, we look forward to performing these experiments for other code models in order to determine whether these findings can extend beyond CodeBERT.
@@ 4 Related Works
Several studies have examined the space of model weights for backdoor attacks and defense. For instance, Garg et al. [32] applied adversarial weight perturbations to inject backdoors into text and image models. Chai and Chen [33] systematically masked the network weights sensitive to backdoors in image models. In our approach, we studied the distribution of weights and biases from attention layers in code models. Among the initial works that utilized activation values is that of Chen et al. [20], which analyzed neuron activations of image models to detect backdoors. They demonstrated that poisoned and benign samples can be grouped into different clusters using these activation values. Additionally, Hussein et al. [34] proposed a technique to create model signatures based on activation optimization, which trains a classifier on the signatures to detect a poisoned model. In our approach, we applied a layer-wise clustering strategy with activation values for code models. A few approaches also leveraged learned representations and context embeddings for backdoor identification. For example, Tran et al. [19] highlighted that poisoned samples might leave unique traces in the feature representations learned by the poisoned model, which can be detectable through spectral signatures. Additionally, Ramakrishnan and Albarghouthi [9], Schuster et al. [15], and Wan et al. [10], among others, adapted the spectral signatures approach for code models. In our approach, we clustered and visualized the context embeddings of the clean and poisoned samples.
5 Conclusion and Future Work
In this paper, we presented a case study for the detection of backdoor signals in CodeBERT, a state-of-the-art large language model of code. We investigated attention weights and biases, activation values, and context embeddings of clean and poisoned models to observe whether there are any potential indications of the presence of trojans. In our experiments with the CodeBERT model for the defect detection task, we observed that the activation values of poisoned samples tend to converge into a single cluster within the upper encoder layers, and the context embeddings of poisoned samples are grouped on the basis of the applied triggers. However, we found no significant differences in attention weights and biases between clean and poisoned models. In the future, we plan to extend our analysis to a broader spectrum of models and tasks.
@@ Appendix A Appendix
In this section, we present additional results from our exploration regarding the parameter analysis of the CodeBERT model in the first encoder layer. As seen in the last encoder layer (sections 3.1 and 3.4), similarly, we noticed no significant deviations in the first encoder layer. We compared layer by layer for the other layers as well; however, we did not obtain any noticeable signal indicating the presence of a backdoor in the poisoned model.

### 05 Mar 2024 v2

@@ 1 Introduction
Large language models (LLMs), trained on extensive publicly available datasets, exhibit remarkable capabilities in software development practices, such as code generation and vulnerability detection ([1]). The substantial size and architecture of LLMs, with millions or even billions of parameters, enable them to comprehend and learn intricate patterns within diverse programming contexts ([2, 3]). These models demonstrate proficiency in handling complex downstream tasks with minimal or no additional fine-tuning, using zero-shot learning or prompting in both natural and formal languages ([4, 5, 6]).
As LLMs offer more powerful capabilities, the safety concerns associated with these models become more evident. Previous studies have demonstrated the susceptibility of code models to adversarial attacks ([7, 8, 9]), the presence of hidden backdoors in code snippets ([10, 11, 12]), or the tendency to memorize data points for decision making ([13, 14, 15, 16]). Specifically, backdoor attacks have become one of the stealthy threats to large language models in recent times, as they are typically compromised with a small amount of poisoned data and manipulate the output of models in malicious ways ([17, 18]). A backdoor attack in the context of coding refers to the insertion of triggers in input programs, allowing attackers to change normal behaviors to erroneous behaviors when exposed to triggers ([19, 20]). Several approaches, such as spectral signatures ([21]), neuron activations ([22]), and occlusion-based approaches ([23]) have been proposed to detect potential poisoned behaviors in large language models of code ([10, 24, 25]).
@@ 2 Proposed Approach
We start with preparing a clean model and a poisoned model, which will serve as the baseline for conducting our experiments. Initially, we load the pre-trained CodeBERT model ([26]) and fine-tune it for the defect detection task ([2]) using the Devign dataset ([27]). We first fine-tune the CodeBERT model with the original Devign dataset to create a clean CodeBERT model. Next, we fine-tune the CodeBERT model separately with the poisoned Devign dataset of variable renaming triggers ([20]) to create a poisoned CodeBERT model. The accuracy of the clean and poisoned models are 63.10% and 62.30%, respectively, and the attack success rate of the latter is 99.22% (see first row of Table 1). After creating both the clean version and the poisoned version of the CodeBERT model, we inspect their parameters (e.g. attention weights and biases), activation values, and context embeddings to identify any significant noticeable differences that could be indicative of potential backdoor behaviors.
@@
Activation Values ([22]). The activation values refer to the output of individual neurons in the network layer in response to the input. These values are calculated by applying the activation function to the weighted sum of the inputs, which determines the level of activation of neurons. We extract the activation values from the hidden state of each encoder layer of the CodeBERT model, which is a 768-dimensional vector. In this experiment, we used all test samples of the Devign dataset.
Context Embeddings ([29]). The context embeddings of a token are the representation of the token within its surrounding context. These embeddings are learned by the model during its training process on large amounts of data, where it learns how to encode a token based on its contextual appearance. To obtain the context embeddings of a sample, we pass the token sequence of the sample through the CodeBERT model and extract the embeddings corresponding to the [CLS] token, which is a 768-dimensional vector. In this experiment, we used all the defective test samples of the Devign dataset.
We adopted the source code provided by Microsoft/CodeXGLUE [2] and Salesforce/CodeT5 [30] to fine-tune the model and extract model parameters. After extracting the attention weights and biases, activation values, and context embeddings from each encoder layer of the pre-trained, clean fine-tuned, and poisoned fine-tuned models, we visualize the distribution of attention weights and biases (see Section 3.1), as well as the clustering of activation values (see Section 3.2) and context embeddings (see Section 3.3). Additionally, we compare the fine-tuned parameters with their corresponding pre-trained parameters (see Section 3.4) and reset the fine-tuned weights to their corresponding pre-trained weights based on a threshold (see Section 3.5). Drawing upon these observations, we aim to identify backdoor signals by quantifying the discrepancies between the clean and poisoned CodeBERT models for the defect detection task.
@@ 3 Experimental Results
3.1 Analysis 1: Distribution of Attention Weight and Bias in Clean and Poisoned Models
@@ 3.3 Analysis 3: Visualizing Context Embeddings of Clean and Poisoned Models
In this analysis, we investigated the applicability of context embeddings in the detection of backdoor signals. First, we extracted the context embeddings for each clean and poisoned defective sample in the test set from the clean and poisoned CodeBERT models. Next, we visualized these context embeddings separately using t-SNE ([33, 32]), as shown in Figure 4. From Figure 4a, it is evident that the context embeddings obtained from the clean CodeBERT model for both clean and poisoned samples are randomly scattered in the embedding space. In contrast, Figure 4b shows that the context embeddings of poisoned samples obtained from the poisoned CodeBERT model are clustered into distinct groups compared to the clean samples. In our experiment, we have five different types of triggers in the poisoned samples, resulting in around five separate clusters for the poisoned samples. Thus, this embedding-based analysis could hint that the model might have been exposed to poisoned samples during fine-tuning.
3.4 Analysis 4: Comparison of Fine-tuned Parameters with Pre-trained Parameters
In this analysis, we chose to represent the differences in peer-to-peer parameters to account for minor parameter fluctuations that might be overlooked in the distribution plots in Figures 1 and 2. Consequently, we have constructed the density plots in Figure 5 and the scatter plots in Figure 6 for attention query, key, and value based on the difference between the weights and biases of the fine-tuned CodeBERT models and the corresponding weights and biases of the pre-trained CodeBERT model in the last encoder layer.
@@ 3.4 Analysis 4: Comparison of Fine-tuned Parameters with Pre-trained Parameters
In Figure 5, we show the differences between the fine-tuned weights and the corresponding pre-trained weights for clean and poisoned CodeBERT models, for each of the three attention components (query, key, and value) of the last encoder layer; we used the Gaussian kernel density estimation ([34]) as the smoothing function [35, 36]. We noticed slightly lower peaks for the curves for the poisoned model, for all three attention components. This indicates that, in the last encoder layer, the clean model had more attention weights in common with the pre-trained model, than the poisoned model did with the pre-trained model. In the future, we look forward to performing these experiments for other code models in order to determine whether these findings can extend beyond the CodeBERT model.
@@
3.5 Analysis 5: Resetting Fine-tuned Weights to Pre-trained Weights based on a Threshold
In this analysis, we conducted an additional experiment to investigate the impact on accuracy (ACC) and attack success rate (ASR) of the poisoned model when resetting its fine-tuned weights ($W_{FT}$) to the corresponding weights from the original pre-trained model ($W_{PT}$). During the fine-tuning process, a model dynamically updates its pre-trained weights for the downstream task; this adaptation might include adjustments toward learning backdoor triggers in the case of a poisoned model. Hence, resetting the weights could serve as a detection mechanism for potential backdoor presence, especially if significant changes in ASR are observed for the poisoned model without a corresponding decrease in ACC. Therefore, we specifically reset the attention weights of the clean and poisoned CodeBERT models that exceeded a predefined threshold in terms of their ratio to the original pre-trained attention weights of the CodeBERT model, such as $\bigl{(}\frac{W_{FT}}{W_{PT}}\sim\text{Threshold}\bigr{)}$. The lower the threshold value is set, the more weight components will be reset. The results of this experiment are presented in Table 1, where we observed a decrease in both ACC and ASR of the model when lowering the threshold value; consequently, this does not point to any indication of backdoor signals.
Table 1: The performance of the clean and poisoned CodeBERT models after resetting their attention weights exceeding a predefined threshold from the original pre-trained CodeBERT model.
Thresholds Clean Model - ACC Poisoned Model - ACC Poisoned Model - ASR
No-Resetting 63.10 62.30 99.22
1.1 62.77 61.64 98.52
1.01 58.09 60.03 85.13
1.001 57.21 60.07 82.78
@@ 4 Related Works
Several studies have examined the space of model weights for backdoor attacks and defense. For instance, Garg et al. [37] applied adversarial weight perturbations to inject backdoors into text and image models. Chai and Chen [38] systematically masked the network weights sensitive to backdoors in image models. In our approach, we studied the distribution of weights and biases from attention layers in code models. Among the initial works that utilized activation values is that of Chen et al. [22], which analyzed neuron activations of image models to detect backdoors. They demonstrated that poisoned and benign samples can be grouped into different clusters using these activation values. Additionally, Hussein et al. [39] proposed a technique to create model signatures based on activation optimization, which trains a classifier on the signatures to detect a poisoned model. In our approach, we applied a layer-wise clustering strategy with activation values for code models. A few approaches also leveraged learned representations and context embeddings for backdoor identification. For example, Tran et al. [21] highlighted that poisoned samples might leave unique traces in the feature representations learned by the poisoned model, which can be detectable through spectral signatures. Additionally, Ramakrishnan and Albarghouthi [10], Schuster et al. [17], and Wan et al. [11], among others, adapted the spectral signatures approach for code models. In our approach, we clustered and visualized the context embeddings of the clean and poisoned samples. Furthermore, Xu et al. [40] and Rajabi et al. [41] trained a binary classifier using outputs from a trojaned and a non-trojaned model as training data to determine whether an unknown model is trojaned or not. In our approach, we used the attention parameters from both clean and poisoned models as training data for the binary classifier for trojaned signal detection.
5 Conclusion
In this paper, we presented a case study for the detection of backdoor signals in CodeBERT, a state-of-the-art large language model of code. We investigated attention weights and biases, activation values, and context embeddings of clean and poisoned models to observe whether there are any potential indications of the presence of trojans. In our experiments with the CodeBERT model for the defect detection task, we observed that the activation values of poisoned samples tend to converge into a single cluster within the upper encoder layers, and the context embeddings of poisoned samples are grouped on the basis of the applied triggers. However, we found no significant differences in attention weights and biases between clean and poisoned models. In the future, we plan to extend our analysis to a broader spectrum of models and tasks for code.
@@ Appendix A Appendix
In this section, we present additional results from our exploration regarding the parameter analysis of the CodeBERT model, including the initial encoder layer (Section A.1), freezing the normalization layers (Section A.2), and the ratio of attention parameters (Section A.3). Additionally, we train a binary classifier using model parameters for identifying clean and poisoned models (see Section A.4).
A.1 Attention Weights and Biases in the Initial Encoder Layer
As observed in the final encoder layer in Sections 3.1 and 3.4, likewise, we observed no significant deviations in the initial encoder layer for attention weights and biases, as depicted in Figures 7, 8, 9 and 10. We compared layer by layer for the other layers as well; however, we did not obtain any noticeable signal indicating the presence of a backdoor in the poisoned model, based on attention weights and biases.
@@ A.1 Attention Weights and Biases in the Initial Encoder Layer
A.2 Clean and Poisoned Attention Parameters after Freezing the Normalization Layers
We also analyzed the weights and biases of the normalization layers, which are typically applied to normalize the output of the attention block, in both the clean CodeBERT model and its poisoned counterpart. Similar to Figure 1 and Figure 2, which illustrate the distribution of attention parameters, we examined the distribution of normalization parameters in different encoder layers. Upon inspecting the parameters of normalization layers, we noticed dissimilarities in the weights and biases between the clean and poisoned models. For instance, Figure 11a illustrates the distribution of the LayerNorm bias for the final encoder layer of the CodeBERT model.
(a) Attention Output Bias (LayerNorm)
(a) Attention Output Bias (LayerNorm)
(b) Attention Query Bias (Freezing)
Figure 11: Comparison of attention biases between the clean (shown as green) and poisoned (shown as red) CodeBERT models in the last encoder layer after freezing the normalization layers.
As a result, we subsequently analyzed the impact of freezing the normalization layers during the fine-tuning process of the CodeBERT model for the defect detection task. Following a similar approach as outlined in Section 3.1, we again generated the distributions of the attention weights and biases for both the clean and poisoned CodeBERT models by freezing their normalization layers. Upon inspecting the attention parameters by freezing the normalization layers, we noticed the CodeBERT model adapts to the defect detection task largely through updating the attention biases compared to the attention weights. For instance, Figure 11b illustrates the comparison of the attention query bias after freezing the normalization layers for the final encoder layer of the CodeBERT model.
A.3 Ratio of Fine-tuned Attention Parameters to Pre-trained Attention Parameters
(a) Attention Weight (Key)
(a) Attention Weight (Key)
(b) Attention Bias (Query)
Figure 12: Visualizing the ratio of fine-tuned attention parameters to pre-trained attention parameters for clean and poisoned CodeBERT models in the last encoder layer.
Similar to the difference in peer-to-peer parameters in Section 3.4, we also represented the ratio in peer-to-peer parameters in this section. In Figure 12a, we constructed the ratio between the attention weights of the poisoned CodeBERT model ($W_{P}$) and the corresponding attention weights of the clean CodeBERT model ($W_{C}$), such as $\frac{W_{P}}{W_{C}}$. To explore the alterations in parameters following the fine-tuning process for both the clean and poisoned CodeBERT models, in Figure 12b, we also represented the ratios of the clean attention weights divided by their corresponding attention weights in the original pre-trained model, and similarly, the ratios of the attention weights in the poisoned model relative to those in the original pre-trained model.
A.4 Training a Binary Classifier for Identifying Clean and Poisoned Models
The distribution of parameters obtained from a poisoned model might be different from that of a non-poisoned clean model [40, 41]. Therefore, we built a binary classifier trained on the attention weights of a set of clean and poisoned models. In particular, we extracted attention weights from CodeBERT models and created training samples by flattening the attention weights into a single dimension. We labeled the attention weights of the poisoned models as 1 and those of the clean models as 0. We then trained a CNN model for the binary classification task to determine whether an unknown model is poisoned or not. However, our trained classifier with code models was unable to distinguish poisoned models from clean models based solely on the model parameters and returned the same label for all test samples. We attempted training different CNN classifiers by adjusting their hyperparameters, yet still observed similar results for the CodeBERT model of the defect detection task.